OpenStudy (anonymous):

factor x^2+5000-6710000 PLzzzzzzz!!!!!!

OpenStudy (tkhunny):

Do you mean "5000x"?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

x^2+5000x-6710000 I might be tempted to Complete the Square since I would guess that 6710000 has far too many factors to sift through. x^2+5000x-6710000 (x^2 + 5000x + _____) - 6710000 - ____ 5000/2 = 2500 2500^2 = 6250000 (x^2 + 5000x + 6250000) - 6710000 - 6250000 (x + 2500)^2 + 12960000 = (x + 2500)^2 - 3600^2 And we see a beautiful Difference of Squares!! (x + 2500)^2 - 3600^2 = ((x+2500)+3600)((x+2500)-3600) = (x + 6100)(x-1100) That was kind of fun!

OpenStudy (anonymous):

thank u very much